OpenStudy (anonymous):

A chemical reaction of the general type A->2B is first-order. with a rate constant of 1.52x10-4s-1. A. Calculate the half-life of A. B.Assuming the initial concentration of A is 0.067mol L-1, calculate the time needed for the concentration to fall to 0.010mol L-1.

OpenStudy (cuanchi):

A. the half life for a first order reaction is independent of the initial concentration of A and respond at the formula t1/2= 0.693/k k= rate constant B. use the integrated equation for the first order reaction ln [A]f = -k t + ln [A]o isolate t = time has to be ~ 2.5 half-lifes